The acceleration of the spacecraft in which the Apollo astronauts took off from the moon was 3.4 m/s2. On the moon, g = 1.6 m/s2. Part A What was the apparent weight of a 75 kg astronaut during takeoff? Express your answer with the appropriate units. Show work.
Answer #1

Given,

a = 3.4 m/s^2 ; g = 1.6 m/s^2

m = 75 kg

Apparent weight will be:

W = (a + g)m

W = 75 (3.4 + 1.6) = 375 N

Hence, W = 375 N